# Pixel-Wise Recognition for Holistic Surgical Scene Understanding

### 20 Jan 2024 v1

@@ 5.2.2 Instrument Segmentation Experiments
Furthermore, we validate our segmentation annotations by training and optimizing ISINet (González et al. (2020)) as a CNN-based baseline and comparing it with our instrument segmentation baseline (called TAPIS Frame), as detailed in Table 11. We evaluate the performance of TAPIS Frame using multiple backbones, including ResNet50, for comparison with ISINet and a Swin Large (SwinL) backbone to leverage the most extensive backbone configuration. Our findings consistently demonstrate the superiority of transformer-based methods, with performance positively correlating with the number of parameters of the backbone. Notably, all instance segmentation methods exceed their detection counterparts in box $mAP@0.5IoU_{box}$, with even the most minor TAPIS Frame configuration achieving comparable performance to the largest detector with less than half of the parameters and FLOPs. These results highlight our segmentation annotations’ impact on training specialized models that capture finer visual cues regarding the pixel-level shape of instruments that improve instrument localization and recognition. Additionally, as shown in Table 11, we test multiple publicly available instrument segmentation methods on our cross-validation set. For this purpose, we optimize and adapt methods designed for instrument segmentation in Endovis 2017 and 2018 by using their official implementations; these methods include TernausNet (Shvets et al. (2018)) [^12] https://github.com/ternaus/robot-surgery-segmentation and MF-TAPNet (Jin et al. (2019)) [^13] https://github.com/keyuncheng/MF-TAPNet for semantic segmentation, and ISINet (González et al. (2020)) [^14] https://github.com/BCV-Uniandes/ISINet and QPD Mask DINO (Dhanakshirur et al. (2023)) [^15] https://github.com/AINeurosurgery/Learnable-QPD-for-maskDINO for Instance Segmentation. These methods yield a relative behavior that aligns with the current state-of-the-art in instrument segmentation and underscores the enhanced performance of our instrument segmentation model over previous approaches. Also, they consistently prove the superiority of Transformer-based architectures.
@@
Moreover, Table 11 also presents the performance of the entire TAPIS model for instrument segmentation (called TAPIS Full), which integrates the region classification head with the TAPIS Frame’s region proposals. The metrics for this experiment, including parameters and FLOPs, focus on the video feature extractor and the region classification head. Including the region classification head enhances performance in the highly stringent semantic segmentation metrics, although with a decrease in $mAP@0.5IoU_{segm}$ performance. We attribute this outcome to the head’s ability to accurately reclassify more prominent and complex regions by correlating them over time. However, it tends to misclassify smaller, independent regions that appear momentarily, resulting in higher recall regarding intersection over union but slightly reduced effectiveness in instance identification. Throughout our experimentation, we noted minimal variance in performance related to the specific design of the time window and the region classification head. However, we observed marginally lower performance with windows covering less than 8 seconds and achieved the optimal performance using a sampling rate of 2fps. Consequently, we opted for a window design comprising 16 frames at a 2fps sampling rate.
@@ 5.2.3 Atomic Action Detection
Additionally, our experiments showed that decreasing the temporal window’s sampling rate per second (increasing the temporal stride) significantly decreases performance, validating our hypothesis about the temporal granularity of the atomic actions detection task. As their name indicates, atomic actions develop in extremely short periods, and they are a highly specific and granular form of modeling surgical activities. During the annotation process, we observed numerous instances where atomic actions unfolded in as few as four consecutive video frames. This observation indicates that the model requires a finer temporal sampling to fully identify actions over time. Consequently, the model performs best with the highest possible sampling rate, allowing for a more detailed and accurate temporal analysis of the surgical actions. This insight is crucial for enhancing the precision and effectiveness of atomic action detection in surgical videos.
@@ 6 Conclusions
Declaration of Generative AI and AI-assisted technologies in the writing process: During the preparation of this work, the author(s) used Grammarly Premium and Open AI’s ChatGPT based on GPT-4 in order to assist in drafting, checking spelling and grammar, organizing text, and paraphrasing or looking for alternative vocabulary. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.
CRediT Author Statement
Nicolás Ayobi:
Conceptualization, Data Curation, Formal Analysis, Investigation, Methodology, Project Administration, Software, Supervision, Validation, Visualization, Writing – original draft, Writing – review & editing.; Santiago Rodríguez: Conceptualization, Data Curation, Formal Analysis, Investigation, Methodology, Software, Validation, Visualization, Writing – original draft.; Alejandra Perez: Conceptualization, Data Curation, Formal Analysis, Investigation, Methodology, Software, Validation, Visualization, Writing – original draft. Isabela Hernandez: Conceptualization, Data Curation, Investigation, Methodology, Project Administration, Software, Supervision, Validation, Writing – review & editing.; Nicolás Aparicio: Data Curation, Formal Analysis, Software, Validation, Visualization, Writing – original draft.; Eugénie Dessevres: Data Curation, Formal Analysis, Software, Validation, Visualization, Writing – original draft.; Sebastián Peña: Data Curation, Resources, Writing – original draft.; Jessica Santander: Data Curation, Resources, Writing – original draft.; Juan Ignacio Caicedo: Data Curation, Project Administration, Resources, Writing – original draft.; Nicolás Fernanez: Conceptualization, Project Administration, Resources, Writing – review & editing.; Pablo Arbelaez: Conceptualization, Funding Acquisition, Methodology, Project Administration, Resources, Supervision, Validation, Writing – review & editing.
@@
Appendix A GraSP General Overview
@@
Appendix B GraSP Statistics
We extend the statistical analysis of frequency distributions across all tasks within GraSP. Tables 161820, and 21 demonstrate that we maintained the consistency in annotation and distribution throughout the two-fold cross-validation setup and in the test set. These tables show the total number of instances per task in each fold, enabling the identification of over-represented and under-represented instances in each task. This analysis provides insights into their correlation with real-life surgical scenarios.
Table 16: Number of instances for each instrument category in the GraSP dataset.
Set Instrument ID Frequency
Fold 1 1 876
2 386
3 524
4 915
5 405
6 28
7 101
Fold 2 1 818
2 355
3 372
4 850
5 411
6 36
7 93
Test 1 809
2 330
3 449
4 844
5 310
6 38
7 81
Table 17: Number of instrument instances in annotated frames.
Set Number of Instances Frequency Percentage
Fold 1 1 52 4.24%
2 511 41.71%
3 492 40.16%
4 165 13.47%
5 5 0.41%
Fold 2 1 65 5.91%
2 412 37.49%
3 446 40.58%
4 172 15.65%
5 4 0.36%
Test 1 67 5.96%
2 500 44.44%
3 442 39.29%
4 112 9.96%
5 4 0.36%
@@ Appendix B GraSP Statistics
Table 18: Number of frames for each class of the atomic action recognition task.
Set Action ID Frames (1fps)
Fold 1 1 93
2 94
3 50
4 43
5 1175
6 48
7 3
8 132
9 242
10 25
11 1883
12 211
13 838
14 5
Fold 2 1 77
2 75
3 30
4 32
5 1013
6 41
7 15
8 99
9 222
10 25
11 1691
12 188
13 813
14 7
Test 1 59
2 71
3 42
4 56
5 923
6 51
7 25
8 108
9 182
10 15
11 1738
12 186
13 737
14 10
@@
Table 19: Atomic actions frequencies and their co-occurrences in the GraSP dataset.
@@
Table 19 presents the co-occurrences between actions in the GraSP dataset. We observe that the most common actions are (hold, still), (still) and (travel). This situation is expected as still and travel are actions fundamental to describing the activities of all instruments and are constantly and widely performed in any surgery instance. The other action shown here is hold. This situation coincides with the action distribution shown in Table 18 where hold is the second most common action. This action cannot be performed alone and, thus, is always combined with another action to describe the activity of an instrument. These action combinations reflect the manipulation of tissues and objects within the surgical scene. We also observe that some sets of actions have a low frequency. These actions correspond to uncommon moves that are nonetheless valid activities that are part of usual procedures and can be seen in other surgeries.
Tables 20 and 21 represent the number of annotations for both tasks. The most frequent phase and step corresponds to idle, and this is because it represents the beginning and end points of the surgery and also transition times between other phases and steps. Figures 20 and 21 respectively show the transitions between phases and steps. These figures show how phase 0 and step 0 are recurrent throughout all surgeries and represent the transition between one step and another and one phase and another. Lastly, both tables and figures show consistency in annotation processes for both steps and phases tasks.
Table 20: Number of frames for each phase in the GraSP dataset.
Set Phase ID Frames (1fps)
Fold 1 0 9938
1 4752
2 3719
3 4367
4 1355
5 3349
6 2899
7 680
8 1752
9 1198
10 4936
Fold 2 0 8192
1 5815
2 3704
3 4506
4 1234
5 2274
6 2159
7 655
8 2529
9 1431
10 2715
Test 0 12934
1 4838
2 3433
3 5317
4 1478
5 2581
6 3608
7 1041
8 3657
9 902
10 3108
Table 21: Number of frames for each step in the GraSP dataset.
Set Step ID Frames (1fps)
Fold 1 0 9917
1 3130
2 2605
3 2060
4 676
5 1612
6 377
7 135
8 2739
9 849
10 698
11 118
12 80
13 718
14 1272
15 992
16 712
17 938
18 5120
19 2895
20 762
Fold 2 0 8192
1 4628
2 2109
3 1578
4 1204
5 1009
6 345
7 1233
8 1893
9 908
10 919
11 35
12 128
13 651
14 1163
15 801
16 559
17 1083
18 4292
19 1916
20 568
Test 0 12934
1 5340
2 3646
3 1366
4 545
5 630
6 481
7 783
8 3223
9 1112
10 456
11 84
12 62
13 818
14 1228
15 1305
16 491
17 1009
18 3506
19 2486
20 1392
The transition matrices in Figures 18 and 19 present the transition probabilities of phases and steps across the GraSP dataset. Each value represents the probability of going from the previous step/phase (rows) to the next step/phase (columns). They demonstrate the logical connections between the phases and the steps. For instance, Figure 18 shows that Developing the space of Retzius always follows the Right pelvic isolated lymphadenectomy, as defined by the surgical team. There can be a pause between the two phases, corresponding to an Idle phase.
@@
The matrix in Figure 22 presents the co-occurrences between the instruments and atomic actions in the GraSP dataset. In other words, this plot shows the typical functions of every instrument in the RARP instances of GraSP. The columns of the plot do not add to 100%, as instruments can perform up to 3 actions simultaneously. The plot reveals that the Bipolar Forceps usually holds tissues or objects, either in a still position or in movement between two locations (travel). The Prograsp Forceps differs from it because of its inability to cauterize. Additionally, it is usually used to hold still bigger tissues in the background passively, which explains why it has the highest realization of still and the second highest of hold. The Large Needle Driver usually stitches tissues with a needle and thread. As such, it is normal for it to be either still or traveling and holding an object. Because the whole purpose of this instrument is to perform the stitching, it travels more than other instruments with pincers that are sometimes used to hold tissues (such as the Bipolar Forceps, the Prograsp Forceps, or the Laparoscopic Instruments). The Monopolar Curved Scissors is the only instrument to be equipped with a tip capable of cutting (apart from a particular type of LI that is observed in one case in the fold 1) and thus has the highest percentage of realization of the action cut. It is also one of the two instruments capable of performing cauterize (the other being the Bipolar Forceps). This action serves to control and prevent bleeding. The Suction Instrument is the only instrument able to perform suction and does it in more than half of the frames in which it is present. The actions it performs in the other action are instances of wash. The Clip Applier is often performing close and release. These actions can be seen when it takes hold or releases the tissue it has to staple. This instrument is not often seen in the dataset and appears for a short amount of time. It usually travels to the tissue, staples it, and leaves the frame. While it is there, no other procedure is performed. This specificity in its use explains why it is the instrument with the highest percentage of travel. Additionally, the action it performs in the other action are instances of staple. Finally, the Laparoscopic Instrument has a similar role to the Prograsp Forceps, presenting a similar distribution. Nonetheless, it is slightly more active and has a higher frequency of the travel action.
Fig. 24: Co-occurrence distribution matrix of instruments in function of the phase in GraSP.
Fig. 25: Co-occurrence distribution matrix of instruments in function of the step in GraSP.
Appendix C TAPIS performance in GraSP

### 26 Jan 2024 v2

@@ 5.2.2 Instrument Segmentation Experiments
Furthermore, we validate our segmentation annotations by training and optimizing ISINet (González et al. (2020)) as a CNN-based baseline and comparing it with our instrument segmentation baseline (called TAPIS Frame), as detailed in Table 12. We evaluate the performance of TAPIS Frame using multiple backbones, including ResNet50, for comparison with ISINet and a Swin Large (SwinL) backbone to leverage the most extensive backbone configuration. Our findings consistently demonstrate the superiority of transformer-based methods, with performance positively correlating with the number of parameters of the backbone. Notably, all instance segmentation methods exceed their detection counterparts in $mAP@0.5IoU_{box}$, with even the most minor TAPIS Frame configuration achieving comparable performance to the largest detector with less than half of the parameters and FLOPs. These results highlight our segmentation annotations’ impact on training specialized models that capture finer visual cues regarding the pixel-level shape of instruments that improve instrument localization and recognition. Additionally, as shown in Table 12, we test multiple publicly available instrument segmentation methods on our cross-validation set. For this purpose, we optimize and adapt methods designed for instrument segmentation in Endovis 2017 and 2018 by using their official public implementations; these methods include TernausNet (Shvets et al. (2018)) [^12] https://github.com/ternaus/robot-surgery-segmentation and MF-TAPNet (Jin et al. (2019)) [^13] https://github.com/keyuncheng/MF-TAPNet for semantic segmentation, and ISINet (González et al. (2020)) [^14] https://github.com/BCV-Uniandes/ISINet and QPD Mask DINO (Dhanakshirur et al. (2023)) [^15] https://github.com/AINeurosurgery/Learnable-QPD-for-maskDINO for Instance Segmentation. These methods yield a relative behavior that aligns with the current state-of-the-art in instrument segmentation and underscores the enhanced performance of our instrument segmentation model over previous approaches. Also, they consistently prove the superiority of Transformer-based architectures.
@@
Moreover, Table 12 also presents the performance of the entire TAPIS model for instrument segmentation (called TAPIS Full), which integrates the region classification head with the TAPIS Frame’s region proposals. The metrics for this experiment, including parameters and FLOPs, focus on the video feature extractor and the region classification head. Including the region classification head enhances performance in the highly stringent semantic segmentation metrics, although with a decrease in $mAP@0.5IoU_{segm}$ performance. We attribute this outcome to the head’s ability to accurately reclassify more prominent and complex regions by correlating them over time. However, it tends to misclassify smaller, independent regions that appear momentarily, resulting in higher recall regarding intersection over union but slightly reduced effectiveness in instance identification. Throughout our experimentation, we noted minimal variance in instrument classification performance related to the specific design of the time window and the region classification head. However, we observed marginally lower performance with windows covering less than 8 seconds and achieved the optimal performance using a sampling rate of 2fps. Consequently, we opted for a window design comprising 16 frames at a 2fps sampling rate.
@@ 5.2.3 Atomic Action Detection
Additionally, our experiments showed that decreasing the temporal window’s sampling rate per second (increasing the temporal stride) significantly decreases performance, hence validating our hypothesis about the temporal granularity of the atomic actions detection task. As their name indicates, atomic actions develop in extremely short periods, and they are a highly specific and granular form of modeling surgical activities. During the annotation process, we observed numerous instances where atomic actions unfolded in as few as four consecutive video frames. This observation indicates that the model requires a finer temporal sampling to fully identify actions over time. Consequently, the model performs best with the highest possible sampling rate, allowing for a more detailed and accurate temporal analysis of the surgical actions. This insight is crucial for enhancing the precision and effectiveness of atomic action detection in surgical videos.
@@ 6 Conclusions
Appendix A Original Frame Pre-Processing
@@ Appendix B GraSP General Overview
Table 1: Number of frames sampled at 1fps for each phase in the GraSP dataset. Phase categories are Idle (0), Left pelvic isolated lymphadenectomy (1), Right pelvic isolated lymphadenectomy (2), Developing the Space of Retzius (3), Ligation of the deep dorsal venous complex (4), Bladder neck identification and transection (5), Seminal vesicle dissection (6), Development of the plane between the prostate and rectum (7), Prostatic pedicle control (8), Severing of the prostate from the urethra (9), and Bladder neck reconstruction (10).
Set Phase Category ID
0 1 2 3 4 5 6 7 8 9 10
Fold 1 9938 4752 3719 4367 1355 3349 2899 680 1752 1198 4936
Fold 2 8192 5815 3704 4506 1234 2274 2159 655 2529 1431 2715
Test 12934 4838 3433 5317 1478 2581 3608 1041 3657 902 3108
Table 2: Number of frames sampled at 1fps for each step in the GraSP dataset. Step categories are Idle (0), Identification and dissection of the Iliac vein and artery (1), Cutting and dissection of the external iliac vein’s lymph node (2), Obturator nerve and vessel path identification, dissection, and cutting of the obturator lymph nodes (3), Insert the lymph nodes in retrieval bags (4), Prevessical dissection (5), Ligation of the dorsal venous complex (6), Prostate dissection until the levator ani (7), Seminal vesicle dissection (8), Dissection of Denonviliers’ fascia (9), Cut the tissue between the prostate and the urethra (10), Hold prostate (11), Insert prostate in retrieval bag (12), Pass suture to the urethra (13), Pass suture to the bladder neck (14), Pull suture (15), Tie suture (16), Suction (17), Cut suture or tissue (18), Cut between the prostate and bladder neck (19), and Vascular pedicle control (20).
Set Step Category ID
0 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20
Fold 1 9917 3130 2605 2060 676 1612 377 135 2739 849 698 118 80 718 1272 992 712 938 5120 2895 765
Fold 2 8192 4628 2109 1578 1204 1009 345 1233 1893 908 919 35 128 651 1163 801 559 1083 4292 1916 568
Test 12934 5340 3646 1366 545 630 481 783 3223 1112 456 84 62 818 1228 1305 491 1009 3506 2486 1392
Appendix C Additional GraSP Statistics
C.1 Frequency Distributions
We extend the statistical analysis of frequency distributions across all tasks within GraSP. Tables 351, and 2 demonstrate that we maintained the consistency in annotation and distribution throughout the two-fold cross-validation setup and in the test set. These tables show the total number of instances per task in each fold, enabling the identification of over-represented and under-represented instances in each task. This analysis provides insights into their correlation with real-life surgical scenarios.
Table 3: Number of segment instances for each instrument category in the GraSP dataset. Instrument categories are Bipolar Forceps (BF), Prograsp Forceps (PF), Large Needle Driver (LND), Monopolar Curved Scissors (MCS), Suction Instrument (SI), Clip Applier (CA), and Laparoscopic Graspers/Instruments (LG).
Set Instrument Category
BF PF LND MCS SI CA LG
Fold 1 876 386 524 915 405 28 101
Fold 2 818 355 372 850 411 36 93
Test 809 330 449 844 310 38 81
@@ C.1 Frequency Distributions
Table 6 presents the frequency distribution of action combinations in the GraSP dataset. We observe that the most common actions are (hold, still), (still) and (travel). This situation is expected as still and travel are actions fundamental to describing the activities of all instruments and are constantly and widely performed in any surgery instance. The other action shown here is hold. This situation coincides with the action distribution shown in Table 5 where hold is the second most common action. This action cannot be performed alone and, thus, is always combined with another action to describe the activity of an instrument. These action combinations reflect the manipulation of tissues and objects within the surgical scene. We also observe that some sets of actions have a low frequency. These actions correspond to uncommon moves that are nonetheless valid activities that are part of usual procedures and can be seen in other surgeries.
Tables 1 and 2 represent the number of annotations for both tasks. The most frequent phase and step corresponds to idle, and this is because it represents the beginning and end points of the surgery and also transition times between other phases and steps.
Table 4: Number of annotated frames containing each of the possible amounts of instrument instances present.
Set Number of Instruments
1 2 3 4 5
Fold 1 51 511 492 165 5
Fold 2 65 412 446 172 4
Test 67 500 442 112 4
@@
Table 5: Number of instances for each class of the atomic action recognition task. Action categories are Cauterize (Cau), Close (Clo), Cut (Cut), Grasp (Gra), Hold (Hol), Open (Open), Open Something (O.Sm), Pull (Pull), Push (Pus), Release (Rel), Still (Sti), Suction (Suc), Travel (Tra) and Other (Oth).
Set Action Category
Cau Clo Cut Gra Hol Ope O.Sm Pul Pus Rel Sti Suc Tra Oth
Fold 1 93 94 50 43 1175 48 3 132 242 25 1883 211 838 5
Fold 2 77 75 30 32 1013 41 15 99 222 25 1691 188 813 7
Test 59 71 42 56 923 51 25 108 182 15 1738 186 737 10
Table 6: Atomic actions combinations frequencies in the GraSP dataset. We present the number of instances of each possible action combination in GraSP.
@@
C.2 Surgical Transitions
Figures 7 and 9 show the transitions between phases and steps. These figures show how phase 0 and step 0 are recurrent throughout all surgeries and represent the transition between one step and another and one phase and another. Lastly, both tables and figures exhibit similarity in the order and duration of each phase and step during the temporal progression of the surgeries, hence showing the intrinsic procedural and ordered nature of surgical procedures and proving the consistency in annotation processes for both steps and phase tasks.
The transition matrices in Figures 5 and 6 present the transition probabilities of phases and steps across the GraSP dataset. Each value represents the probability of going from the previous step/phase (rows) to the next step/phase (columns). They demonstrate the logical connections between the phases and the steps. For instance, Figure 5 shows that Developing the space of Retzius always follows the Right pelvic isolated lymphadenectomy, as defined by the surgical team. There can be a pause between the two phases, corresponding to an Idle phase. The Figures show that most phases and steps transition from or to an Idle and that steps present more variability in their transition probabilities as step categories can show repeatedly throughout the procedure. Additionally, Figures 8 and 10 show the complete boxplots of the distribution of the duration of each phase and step categories. Besides the previous analysis in the Dataset Statistics Section, we observe that phase categories present more outliers, tending to have more extensive durations, once again proving the temporal granularity difference between phases and steps, where phases tend to represent broader temporal segments.
C.3 Co-Occurence Matrices
Figure 11 portrays the theoretical matrixes established by the experienced surgical team. These matrices demonstrate possible combinations of actions that can occur simultaneously in the same instrument instance, and also the possible actions that each instrument can perform. We design the matrices to have a predefined a priori rule of possible annotations based on the extensive knowledge of our surgeons. We applied this matrix after the first annotation rounds to validate and correct our annotations.
Fig. 12: Co-occurrence distribution matrix of instruments in function of the surgical phase in GraSP.
Fig. 13: Co-occurrence distribution matrix of instruments in function of the surgical step in GraSP.
@@ C.3 Co-Occurence Matrices
Figures 12 and 13 display the inter-task co-occurrence matrices for instruments-phases and instruments-steps. These matrices demonstrate the varying frequency of distinct instruments during different stages of the surgery. Intuitively, the Large Needle Driver (LND) has the highest co-occurrence with the reconstructive phases of the surgeries like the Ligation of the dorsal venous complex and Bladder neck reconstruction phases, which involve the use of the LND to maneuver the surgical needles and threats to suture, ligate and reconstruct tissue. Hence, the LND is also the most frequent in suturing steps like Pass suture to urethra or bladder neck and pull or tie suture. Similarly, the Bipolar Forceps (BF) and Monopolar Curved Scissors are present in most phases and steps but especially towards phases like Left or Right pelvic isolated lymphadenectomy, Bladder neck identification and transection, and in steps like Identification and dissection of the Iliac vein and artery and Seminal vesicle dissection, which are mostly related to tissue manipulation and cutting. Moreover, the Clip Applier (CA) has a very high co-occurrence with the Prostatic pedicle control phase and the Vascular pedicle control step, which require tissue clipping for bleeding control. Also, the Laparoscopic Graspers/Instruments (LG or LI) show mostly in phases and steps requiring the holding of tissue or retrieval bags. Thus, these matrices again demonstrate the strong correlations between different semantic levels of surgical scene understanding.
@@
The matrix in Figure 14 presents the co-occurrences between the instruments and atomic actions in the GraSP dataset. In other words, this plot shows the typical functions of every instrument in the RARP instances of GraSP. The columns of the plot do not add to 100%, as instruments can perform up to 3 actions simultaneously. The plot reveals that the Bipolar Forceps usually holds tissues or objects, either in a still position or in movement between two locations (travel). The Prograsp Forceps differs from it because of its inability to cauterize. Additionally, it is usually used to hold still bigger tissues in the background passively, which explains why it has the highest realization of still and the second highest of hold. The Large Needle Driver usually stitches tissues with a needle and thread. As such, it is normal for it to be either still or traveling and holding an object. Because the whole purpose of this instrument is to perform the stitching, it travels more than other instruments with pincers that are sometimes used to hold tissues (such as the Bipolar Forceps, the Prograsp Forceps, or the Laparoscopic Graspers/Instruments). The Monopolar Curved Scissors is the only instrument to be equipped with a tip capable of cutting (apart from a particular type of LI that is observed in one case in the fold 1) and thus has the highest percentage of realization of the action cut. It is also one of the two instruments capable of performing cauterize (the other being the Bipolar Forceps). This action serves to control and prevent bleeding. The Suction Instrument is the only instrument able to perform suction and does it in more than half of the frames in which it is present. The actions it performs in the other action are instances of wash. The Clip Applier is often performing close and release. These actions can be seen when it takes hold or releases the tissue it has to staple. This instrument is not often seen in the dataset and appears for a short amount of time. It usually travels to the tissue, staples it, and leaves the frame. While it is there, no other procedure is performed. This specificity in its use explains why it is the instrument with the highest percentage of travel. Additionally, the action it performs in the other action are instances of staple. Finally, the Laparoscopic Instrument has a similar role to the Prograsp Forceps, presenting a similar distribution. Nonetheless, it is slightly more active and has a higher frequency of the travel action.
Appendix D TAPIS performance in GraSP